OpenStudy (anonymous):

What is the sum of the arithmetic sequence 152, 138, 124, , if there are 24 terms? (1 points) -240 -216 -192 -168

hartnn (hartnn):

could you find the common difference ?

hartnn (hartnn):

the formula is \(\Large S_n = (n/2) (2a_1 +(n-1)d)\) a1 = 1st term = 152 n = number of terms = 24 you just need d= common difference = ... to be able to use this formula

hartnn (hartnn):

d= common difference = difference between the consecutive terms d = 2nd term - 1st term = 3rd term - 2nd term = .... and so on

OpenStudy (anonymous):

Ok Let Me Try To Solve It

hartnn (hartnn):

first find 'd'

OpenStudy (anonymous):

(24/2) \[\left( \frac{ 24 }{ 2} \right)\left( 2\times152 \right)+\left( 24-1 \right)d\]

hartnn (hartnn):

first find 'd' !! d = 2nd term - 1st term = 3rd term - 2nd term = ....

OpenStudy (anonymous):

Im Kind of confused :/

OpenStudy (anonymous):

About d

hartnn (hartnn):